# DEIM: DETR with Improved Matching for Fast Convergence

### v1

@@ Abstract
We introduce DEIM, an innovative and efficient training framework designed to accelerate convergence in real-time object detection with Transformer-based architectures (DETR). To mitigate the sparse supervision inherent in one-to-one (O2O) matching in DETR models, DEIM employs a Dense O2O matching strategy. This approach increases the number of positive samples per image by incorporating additional targets, using standard data augmentation techniques. While Dense O2O matching speeds up convergence, it also introduces numerous low-quality matches that could affect performance. To address this, we propose the Matchability-Aware Loss (MAL), a novel loss function that optimizes matches across various quality levels, enhancing the effectiveness of Dense O2O. Extensive experiments on the COCO dataset validate the efficacy of DEIM. When integrated with RT-DETR and D-FINE, it consistently boosts performance while reducing training time by 50%. Notably, paired with RT-DETRv2, DEIM achieves 53.2% AP in a single day of training on an NVIDIA 4090 GPU. Additionally, DEIM-trained real-time models outperform leading real-time object detectors, with DEIM-D-FINE-L and DEIM-D-FINE-X achieving 54.7% and 56.5% AP at 124 and 78 FPS on an NVIDIA T4 GPU, respectively, without the need for additional data. We believe DEIM sets a new baseline for advancements in real-time object detection. Our code and pre-trained models are available at https://github.com/ShihuaHuang95/DEIM.
@@ 1 Introduction
However, slow convergence remains one of the primary limitations of DETRs, and we hypothesize that the reasons are two-fold. ❶ Sparse supervision: The O2O matching mechanism assigns only one positive sample per target, greatly limiting the number of positive samples. In contrast, O2M generates several times more positive samples. This scarcity of positive samples restricts dense supervision, which impedes effective model learning—particularly for small objects, where dense supervision is crucial for performance. ❷ Low-quality matches: Unlike traditional methods that rely on dense anchors (usually >8000>8000), DETR employs a small number (100 or 300) of randomly initialized queries. These queries lack spatial alignment with targets, leading to numerous low-quality matches in the training, where matched boxes have low IoU with the targets but high confidence scores.
@@ 2 Related Work
Object detection with transformer (DETR) [3] represents a shift from traditional CNN architectures to transformers. By using Hungarian [16] loss for one-to-one matching, DETR eliminates the need for hand-crafted NMS as the post-processing and enables end-to-end object detection. However, it suffers from slow convergence and dense computation.
@@ 3.1 Preliminaries
Focal loss (FL) [19] was introduced to prevent an abundance of easy negatives from overwhelming the detector during training, directing focus instead towards a sparse set of hard examples. It serves as the default classification loss in DETR-based detectors [45, 39] and is defined as follows:
@@
Figure 2: Anchor/Query Match Comparison. Comparison of the number of matched anchors/queries per image in one $CO$ epoch using one-to-many (SimOTA [44]) and one-to-one (Hungarian [3]) matching schemes.
@@ 3.3 Improving matching quality: Matchability-Aware Loss
These issues are less problematic for traditional detectors due to their dense anchors and multiple assignment strategies. However, in the DETR framework, where queries are sparse and matching is more rigid, these limitations become more pronounced.
@@
4 Experiments
4.1 Training details
Table 1: Comparison with real-time object detectors on COCO [20] val2017. By integrating our method into D-FINE-L [27] and D-FINE-X [27], we build DEIM-D-FINE-L and DEIM-D-FINE-X. We compare our method with YOLO-based and DETR-based real-time object detectors. $\star$} indicates that the NMS is tuned with a confidence threshold of 0.01.
@@ 4.2 Comparisons with real-time detectors
When compared directly to YOLOv11-X [13], our method outperforms this state-of-the-art YOLO model, achieving slightly higher performance (54.7 vs. 54.1 AP) and reducing inference time by 20% (8.07 ms vs. 10.74 ms). Although YOLOv10 [34] uses a hybrid O2M and O2O assignment strategy, our models consistently outperform YOLOv10, demonstrating the effectiveness of our Dense O2O strategy.
@@ 4.3 Comparisons with ResNet [14]-based DETRs
4.4 Comparisons on CrowdHuman
CrowdHuman [30] is a benchmark dataset designed to evaluate object detectors in dense crowd scenarios. We applied both D-FINE and our proposed method to the CrowdHuman dataset, following the configurations provided in the official repository  [^1] https://github.com/Peterande/D-FINE/configs/dfine/crowdhuman. As shown in Tab. 3, our approach (D-FINE-L enhanced with DEIM) achieves a notable improvement of 1.5 AP over D-FINE-L. In particular, our method delivers a significant performance boost (greater than 3% improvement) on small objects (APs) and high-quality detections (AP75), demonstrating its ability to detect objects more accurately in challenging scenarios. Furthermore, this experiment underscores the strong generalization capability of our method across diverse datasets, further confirming its robustness.
@@ 4.4 Comparisons on CrowdHuman
4.5 Analysis
Methods for achieving Dense O2O.
We explored two approaches to implement Dense O2O: mosaic [1] and mixup [38]. Mosaic is a data augmentation technique that combines four images into one, while mixup overlays two images at a random ratio. Both methods effectively increase the number of targets per image, enhancing supervision during training.
As shown in Tab. 4, both mosaic and mixup lead to significant improvements after 12 epochs compared to training without target augmentation, highlighting the effectiveness of Dense O2O. Moreover, combining mosaic and mixup accelerates model convergence, further emphasizing the benefits of augmented supervision. We further tracked the number of positive samples per image over one training epoch, with results shown in Fig. 5. Compared to traditional O2O matching, Dense O2O significantly increases the number of positive samples.
@@ 5 Conclusion
Acknowledgements. Thanks to Xuanlong Yu, Longfei Liu, and Haiyang Xie from Intellindust for their kind discussions and helpful suggestions.
@@ 7 Comparison with Lighter YOLO Detectors
Table 9: Comparison with S and M sized real-time object detectors on COCO [20] val2017. $\star$} indicates that the NMS is tuned with a confidence threshold of 0.01.
@@ 8 Additional Results
Effectiveness of the minor modifications. We incorporate minor modifications, including unfreezing the BN layers in the Backbone, adopting the FlatCosine LR scheduler, and replacing the Decoder activation function with SiLU, into both D-FINE-L and D-FINE-X. After training for 36 epochs, we observe that these changes have no impact on D-FINE-L but lead to a 0.1 AP improvement for D-FINE-X (55.4 vs. 55.5). This configuration is used as the new baseline for our experiments.
Number of positive samples between with/without Dense O2O. During one epoch of training, we compared the number of positive samples in the same training images with and without using Dense O2O, as shown in Fig. 5. After incorporating Dense O2O, the number of positive samples significantly increases. This further supports our claim that Dense O2O effectively enhances supervision.
Figure 5: #\# Positive Samples with and without Dense O2O in One Epoch of Training. Base indicates without Dense O2O.
Figure 6: Qualitative Comparison between D-FINE-L and DEIM. In each paired image, the left is from D-FINE-L while the right is predicted by DEIM-D-FINE-L (Score threshold == 0.5).

### v2

@@ Abstract
We introduce DEIM, an innovative and efficient training framework designed to accelerate convergence in real-time object detection with Transformer-based architectures (DETR). To mitigate the sparse supervision inherent in one-to-one (O2O) matching in DETR models, DEIM employs a Dense O2O matching strategy. This approach increases the number of positive samples per image by incorporating additional targets, using standard data augmentation techniques. While Dense O2O matching speeds up convergence, it also introduces numerous low-quality matches that could affect performance. To address this, we propose the Matchability-Aware Loss (MAL), a novel loss function that optimizes matches across various quality levels, enhancing the effectiveness of Dense O2O. Extensive experiments on the COCO dataset validate the efficacy of DEIM. When integrated with RT-DETR and D-FINE, it consistently boosts performance while reducing training time by 50%. Notably, paired with RT-DETRv2, DEIM achieves 53.2% AP in a single day of training on an NVIDIA 4090 GPU. Additionally, DEIM-trained real-time models outperform leading real-time object detectors, with DEIM-D-FINE-L and DEIM-D-FINE-X achieving 54.7% and 56.5% AP at 124 and 78 FPS on an NVIDIA T4 GPU, respectively, without the need for additional data. We believe DEIM sets a new baseline for advancements in real-time object detection. Our code and pre-trained models are available at https://www.shihuahuang.cn/DEIM/.
@@ 1 Introduction
However, slow convergence remains one of the primary limitations of DETRs, and we hypothesize that the reasons are two-fold. ❶ Sparse supervision: The O2O matching mechanism assigns only one positive sample per target, greatly limiting the number of positive samples. In contrast, O2M generates several times more positive samples. This scarcity of positive samples restricts dense supervision, which impedes effective model learning—particularly for small objects, where dense supervision is crucial for performance. ❷ Low-quality matches: Unlike traditional methods that rely on dense anchors (usually $>8000$), DETR employs a small number (100 or 300) of randomly initialized queries. These queries lack spatial alignment with targets, leading to numerous low-quality matches in the training, where matched boxes have low IoU with the targets but high confidence scores.
@@ 2 Related Work
Object detection with transformer (DETR) [3] represents a shift from traditional CNN architectures to transformers. By using Hungarian [16] algorithm for one-to-one matching, DETR eliminates the need for hand-crafted NMS as the post-processing and enables end-to-end object detection. However, it suffers from slow convergence and dense computation.
@@ 3.1 Preliminaries
Focal loss (FL) [19] was introduced to prevent an abundance of easy negatives from overwhelming the detector during training, directing focus instead towards a sparse set of hard examples. It serves as the default classification loss in DETRs [45, 39] and is defined as follows:
@@
Figure 2: Anchor/Query Match Comparison. Comparison of the number of matched anchors/queries per image in one COCO epoch using one-to-many (SimOTA [44]) and one-to-one (Hungarian [3]) matching schemes.
@@ 3.3 Improving matching quality: Matchability-Aware Loss
These issues are less problematic for traditional detectors due to their dense anchors and one-to-many assignment strategies. However, in the DETR framework, where queries are sparse and matching is more rigid, these limitations become more pronounced.
@@
[…]
@@ 4.2 Comparisons with real-time detectors
When compared directly to YOLOv11-X [13], our DEIM-D-FINE-L outperforms this SoTA model, achieving slightly higher performance (54.7 vs. 54.1 AP) and reducing inference time by 20% (8.07 ms vs. 10.74 ms). Although YOLOv10 [34] uses a hybrid O2M and O2O assignment strategy, our models consistently outperform YOLOv10, demonstrating the effectiveness of Dense O2O.
@@ 4.3 Comparisons with ResNet [14]-based DETRs
4.4 Comparisons on CrowdHuman
CrowdHuman [30] is a benchmark dataset designed to evaluate object detectors in dense crowd scenarios. We applied both D-FINE and our proposed method to the CrowdHuman dataset, following the configurations provided in the D-FINE. As shown in Tab. 3, our approach (D-FINE-L enhanced with DEIM) achieves a notable improvement of 1.5 AP over D-FINE-L. In particular, our method delivers a significant performance boost (greater than 3% improvement) on small objects (APs) and high-quality detections (AP75), demonstrating its ability to detect objects more accurately in challenging scenarios. Furthermore, this experiment underscores the strong generalization capability of DEIM across diverse datasets, confirming its robustness.
@@ 4.4 Comparisons on CrowdHuman
4.5 Analysis
Methods for achieving Dense O2O.
We explored two approaches to implement Dense O2O: mosaic [1] and mixup [38]. Mosaic is a data augmentation that combines four images into one, while mixup overlays two images at a random ratio. Both methods effectively increase the number of targets per image, enhancing supervision during training.
As shown in Tab. 4, both mosaic and mixup lead to significant improvements after 12 epochs compared to training without target augmentation, highlighting the effectiveness of Dense O2O. Moreover, combining mosaic and mixup accelerates model convergence, further emphasizing the benefits of augmented supervision. We further tracked the number of positive samples per image over one training epoch, with results shown in Fig. 5. Compared to traditional O2O, Dense O2O significantly increases positive samples.
@@ 4.5 Analysis
Training speed.
We provide an efficient implementation using Mosaic with caching and Mixup within batches. Tab. 7 shows the one-epoch training time on a single 4090 GPU, where DEIM is almost as fast as the baseline (1.183 vs. 1.181) and requires less training time to converge (71 vs. 85 hours). This highlights that our approach improves convergence while maintaining efficiency.
Table 7: Training time in GPU hours.
Method Epoch #GPU hr AP
RT-DETRv2-R50 1 1.181 -
w/ DEIM 1 1.183 -
[…]
Finetuning from Object365.
We directly employ the pre-trained Object365 weights from D-FINE and compare the results of fine-tuning with and without DEIM. As shown in Tab. 8, DEIM achieves better performance with fewer fine-tuning epochs. It further validates that DEIM delivers consistent gains, even when pre-trained on larger datasets.
Table 8: Fine-tuned results from Object365 pre-training.
Method Epoch AP AP50 AP75
D-FINE-X 32 59.3 76.8 64.6
w/ DEIM 24 59.5 76.4 65.2
@@ 5 Conclusion
Acknowledgements. We thank Xuanlong Yu, Longfei Liu, and Haiyang Xie for their valuable discussions. This work was supported by the Horizontal Project of Hefei Normal University (No. HXXM2022236) and the Key Projects of the National Natural Science Foundation of Universities in Anhui Province (No. 2023AH051302).
@@ 2 Comparison with Lighter YOLO Detectors
[…]
@@ 3 Additional Results
Effectiveness of the minor modifications.
We incorporate minor modifications, including unfreezing the BN layers in the Backbone, adopting the FlatCosine LR scheduler, and replacing the Decoder activation function with SiLU, into both D-FINE-L and D-FINE-X. After training for 36 epochs, we observe that these changes have no impact on D-FINE-L but lead to a 0.1 AP improvement for D-FINE-X (55.4 vs. 55.5). This configuration is used as the new baseline for our experiments.
Number of positive samples between with/without Dense O2O.
During one epoch of training, we compared the number of positive samples in the same training images with and without using Dense O2O, as shown in Fig. 5. After incorporating Dense O2O, the number of positive samples significantly increases. This further supports our claim that Dense O2O effectively enhances supervision.
Figure 5: ##\## Positive Samples with and without Dense O2O in One Epoch of Training. Base indicates without Dense O2O.
Table 12: Varying the number of objects per training image.
Avg ##\## objects AP AP50 AP75
Training 24 Epochs
$\sim$ 10 51.7 69.5 55.8
$\sim$ 25 52.5 70.6 56.7
$\sim$ 50 52.2 70.1 56.4
Table 13: Training and validation accuracy.
Model APtrain APval
RT-DETRv2-R50 65.1 53.4
w/ DEIM 64.8 54.3
Studies of the number of positive samples.
We adjust the average number of objects per image during training by modifying Dense O2O. As shown in Tab. 12, performance improves significantly when the number increases from 10 (without Dense O2O) to 25 (Default Dense O2O) but drops at 50 (Max Dense O2O). This decline is likely due to an imbalance in the positive-to-negative ratio and a data distribution shift caused by too many objects. Notably, an average of 25 objects aligns with the default experimental setting used in this study, corresponding to the default Dense O2O configuration.
Training vs. validation accuracy.
As shown in Tab. 13, DEIM achieves higher validation accuracy and slightly lower training accuracy, indicating reduced overfitting on the training set and improved adaptability to new samples.
Figure 6: Qualitative Comparison between D-FINE-L and DEIM. In each paired image, the left is from D-FINE-L while the right is predicted by DEIM-D-FINE-L (Score threshold $=$ 0.5).